# Current-induced near-field radiative energy, linear-momentum, and angular-momentum transfer

### 13 Dec 2023 v1

@@ Abstract
In this work, we study the near-field radiative energy, linear-momentum, and angular-momentum transfer from a current-biased graphene to nanoparticles. The electric current through the graphene sheet induces nonequilibrium fluctuations, causing energy and momentum transfer even in the absence of a temperature difference. The inherent spin-momentum locking of graphene surface plasmon polaritons leads to an in-plane torque perpendicular to the electric current. In the presence of a temperature difference, energy transfer is greatly enhanced while the lateral force and torque remains within the same order. Our work explores the potential of utilizing current-biased graphene to manipulate nanoparticles.
@@ I Introduction
Nonreciprocity-induced lateral force has been studied recently Gelbwaser-Klimovsky et al. (2021); Khandekar et al. (2021); Tsurimaki et al. (2023); Shapiro (2017). It has been found that the equilibrium lateral force in nonreciprocal systems can be used for propulsion Gelbwaser-Klimovsky et al. (2021). Furthermore, it has been discovered Khandekar et al. (2021) that when at least one of the interacting surfaces is made of nonreciprocal materials, materials with different temperatures can experience a nonequilibrium lateral force in the direction parallel to the surfaces. In the case where some part of the system is affected by mobile carrier drift, an atom or nanoparticle placed above the sample with mobile carrier drift experiences a fluctuation force Shapiro (2017).
@@
Current-biased graphene is also known for its nonreciprocity Borgnia et al. (2015); Duppen et al. (2016); Morgado and Silveirinha (2018, 2020). The exceptional electron mobility of graphene allows it to sustain nonreciprocal surface plasmon polaritons when driven by a current with drift velocity, which gives rise to interesting properties such as Fizeau drag Dong et al. (2021); Zhao et al. (2021) and negative Landau damping Morgado and Silveirinha (2017). Current-biased graphene is in a nonequilibrium state and this leads to a finite photonic chemical potential Henry and Kazarinov (1996) that depends on the in-plane wave vector for the thermal electromagnetic radiation. The occupation number of radiative photons is nonreciprocal, that is asymmetric with respect to the positive and negative in-plane wave vectors Volokitin and Persson (2008, 2011a, 2011b); Peng and Wang . The nonreciprocal properties of current-biased graphene can be adjusted through chemical doping or gate voltage, making it suitable for a wide range of applications in controlling thermal radiation. For example, near-field radiative heat transfer has been studied between two suspended graphene sheets Zhang et al. (2020a), graphene covered substrates Shi et al. (2022), graphene-nanoparticles Zhang et al. (2020b), and graphene-based multilayer systems Zhou et al. (2020). However, the physical aspects of the force and torque experience by the nanoparticle near the current-biased graphene remain unexplored to date.
@@
In this work, we investigate the near-field transfer of energy, linear momentum, and angular momentum between an isotropic dipolar nanoparticle and the current-biased graphene. The nonequilibrium state in the current-biased graphene induces energy transfer between the particle and the graphene even in the absence of a temperature difference. Importantly, these SPPs carry both linear and spin-locked angular momentum. As a result, there is an exchange of net linear and spin angular momentum between the particle and the SPPs, leading to both lateral force and torque experienced by the nanoparticle.
@@
Figure 1: (a) Schematic plot of the lateral force and torque exerted on an isotropic dipolar particle above a current-biased graphene in the near-field regime. The drift velocity $v_{d}$ is along the $x$ direction. (b) Dispersion of the surface plasmon polaritons of the current-biased graphene. At large drift velocities, the drift current effectively drags the SPPs, leading to the nonreciprocal propagation of SPPs. (c) The absolute value of photon occupation number difference $|\delta n(\omega,q_{x})|$ versus $\omega$ with $v_{d}=0.7v_{F}$. The solid and dashed lines are for positive $\delta n(\omega,q_{x})$ for $q_{x}{>}0$ and negative $\delta n(\omega,q_{x})$ for $q_{x}{<}0$, respectively.
@@ II System and formalism
The system is illustrated schematically in Fig. 1(a), where the spherical nanoparticle with radius $R$ is placed near the surface of current-biased monolayer graphene. Without losing generality, we assume that the drift velocity $v_{d}$ in graphene is along the positive direction of the $x$ axis. Within the framework of fluctuational electrodynamics, the total power $H$, lateral force along $x$-direction $F_{x}$, and torque along $y$-direction $M_{y}$ can be expressed as
$P=\int_{0}^{\infty}\frac{d\omega}{2\pi}p(\omega)$ (1)
with $P\in\{H,F_{x},M_{y}\}$ and $p\in\{h,f_{x},m_{y}\}$. The spectral densities are given by Khandekar et al. (2021):
$h(\omega)$ $=-\hbar\omega k_{0}^{2}{\rm Im}[\alpha(\omega)]\int\frac{d^{2}{\bf q}}{4\pi^{2}}\Theta(\omega,{\bf q})\delta n(\omega,q_{x}),$ (2)
$f_{x}(\omega)$ $=\hbar k_{0}^{2}{\rm Im}[\alpha(\omega)]\int\frac{d^{2}{\bf q}}{4\pi^{2}}q_{x}\Theta(\omega,{\bf q})\delta n(\omega,q_{x}),$ (3)
$m_{y}(\omega)$ $=-\hbar{\rm Im}[\alpha(\omega)]\int\frac{d^{2}{\bf q}}{4\pi^{2}}q_{x}{\rm Im}\big{(}r_{p}e^{2i\beta_{0}d}\big{)}\delta n(\omega,q_{x}),$ (4)
@@
$\Theta(\omega,{\bf q})={\rm Re}\bigg{[}\frac{r_{p}e^{2i\beta_{0}d}}{2\beta_{0}}\left(\frac{2q^{2}}{k_{0}^{2}}-1\right)\bigg{]}.$ (5)
Here, $r_{p}$ is the Fresnel reflection coefficient for $p$ polarization. The in-plane wave vector and the angular frequency are denoted by ${\bf q}=(q_{x},q_{y})$ and $\omega$. The magnitude of the out-of-plane wave vector in air is $\beta_{0}=\sqrt{k_{0}^{2}-q^{2}}$ with $k_{0}=\omega/c$ and $q=|{\bf q}|$. The photon-occupation difference between the nanoparticle and the graphene is Volokitin and Persson (2008, 2011a, 2011b)
@@
where $T_{e}$ and $T_{p}$ represent the temperatures of graphene and the nanoparticle, respectively. Equation  (6) indicates that the electric current-induced nonequilibrium fluctuations can produce a finite energy flux even in the absence of a temperature difference. In the following, we assume that the temperature of the particle and the graphene is the same with $T_{p}=T_{e}=T=300\,$K, unless there is a special account for the temperatures. The energy transmission function of the energy flux in Eq.(2) is defined as:
${\cal Z}\left(\omega,{\bf q}\right)=-\frac{1}{4\pi^{2}}\hbar\omega k_{0}^{2}{\rm Im}[\alpha(\omega)]\Theta(\omega,{\bf q})\delta n(\omega,q_{x}),$ (7)
which gives the energy transfer at given $\omega$ and $\bf q$.
@@
The dispersion relation of SPPs in a monolayer graphene can be obtained through the following relation Gonçalves and Peres (2016):
$2i\omega\varepsilon_{0}=\beta_{0}\sigma_{g},$ (8)
where the surface conductivity $\sigma_{g}$ is given by Svintsov and Ryzhii (2019)
@@
with $\mu(T)=2k_{B}T\ln[2\cosh(\mu_{g}/2k_{B}T)]$. Here, $T$ is the temperature, $\mu_{g}$ is the chemical potential, $\gamma$ the damping parameter, and $e_{0}$ the electron charge. In the numerical calculation, we set the Fermi velocity $v_{F}=10^{6}\,$m/s, the chemical potential $\mu_{g}=0.1\,$eV and the damping parameter $\gamma=3.7\,$meV. Figure 1(b) shows the nonreciprocal dispersions of SPPs in the presence of an electric current.
@@ III Numerical Results
Before considering nonequilibrium energy and momentum transfer, we first discuss the nonreciprocal properties of $\delta n(\omega,q_{x})$ shown in Fig. 1(c). We observe that the photon occupation number difference $\delta n(\omega,q_{x})$ is positive for positive $q_{x}$ under $\omega{>}q_{x}v_{d}$. This indicates that the effective photonic temperature of the graphene is higher than the particle temperature $T_{p}$. Furthermore, it should be noted that contributions from $\omega{<}q_{x}v_{d}$ can be disregarded. This is because these contributions become significant only when the separation between the particle and the graphene reaches subnanometer scales Volokitin and Persson (2011b), whereas in our case, the distance is on the order of micrometers. Consequently, photons, especially those with positive $q_{x}$, flow from the graphene to the particle. In contrast, for negative $q_{x}$, $\delta n(\omega,q_{x})$ is always negative. Therefore, the effective photonic temperature of the graphene is lower than $T_{p}$. This sign reversal indicates a reversal in the direction of photon transfer compared to the behavior exhibited by positive $q_{x}$.
We analyze three kinds of nanoparticle: NaCl, undoped GaAs, and InSb. The intricate dielectric characteristics are given by the well-established Lorentz-Drude model with Foteinopoulou et al. (2019); Khandekar and Jacob (2019)
@@
where $\varepsilon_{\infty}$ is the high-frequency dielectric constant, $\omega_{L}$ is the longitudinal optic-phonon frequency, $\omega_{T}$ is the transverse optic-phonon frequency, $\Gamma$ is the optic-phonon damping constant. Within the electrostatic limit, we can reasonably neglect multiple scattering interactions between the particle and the graphene, and describe the particle effectively using its electrical polarizability which is expressed as $\alpha(\omega)=4\pi R^{3}[\varepsilon(\omega)-1)]/[\varepsilon(\omega)+2]$ Zhao et al. (2012). It is important to note that these assumptions are valid only under specific conditions: $R\ll d\ll\hbar c/k_{B}T$. In Fig. 2(a), we present the imaginary part of the polarizability for these nanoparticles, revealing distinct dipolar resonances.
Now we study the effect of drift velocity on the net power transfer experienced by different nanoparticles, as shown in Fig. 2(b). We differentiate between positive and negative energy fluxes, which represent the flow of energy from the graphene to the particle and from the particle to the graphene, respectively. Our analysis reveals intriguing behaviors. For NaCl and InSb particles, the power transfer initially increases as $v_{d}$ increases. However, with increasing $v_{d}$ further, the power transfer decreases. In contrast, the direction of net power transfer undergoes a change for GaAs particles as the drift velocity increases. These are due to the interplay between the nonreciprocal properties of the SPPs and photonic occupation number of graphene.
@@
To analyze the contributions of forward ($q_{x}{>}0$) and backward ($q_{x}{<}0$) SPPs to net power transfer, we use the NaCl particle as an example. In Fig. 2(c), we show the energy transmission function and the dispersion of graphene plasmons (represented by the cyan dashed line). This visual representation aligns the bright branches of the energy transmission function with the cyan dashed line, confirming the dominant role played by SPPs in governing power transfer in our system. For $q_{x}{>}0$, the energy transfer coefficients are positive, indicating energy transfer from graphene to the particle. For $q_{x}{<}0$, the energy transfer coefficients are negative, signifying energy transfer from the particle back to the graphene. Therefore, the net energy transfer is the sum of contributions from $q_{x}{>}0$ and $q_{x}{<}0$. At lower drift velocities, the nonreciprocal characteristics of SPPs are not pronounced, and energy transfer is mainly governed by the nonreciprocal photon occupation number difference $\delta n(\omega,q_{x})$, as shown in Fig. 1(c). However, as the drift velocity increases, nonreciprocal SPPs gradually interact with $\delta n(\omega,q_{x})$. The high drift velocity suppresses the net power transfer caused by the nonreciprocal photon occupation number, resulting in a decrease in energy transfer between NaCl, InSb, and graphene, and shifts in energy transfer direction observed between GaAs and graphene.
@@
Subsequently, we investigate the effect of drift velocity on linear-momentum and angular-momentum transfer between particles the graphene. We define positive $M_{y}$ as directed along the negative $y$-axis and positive $F_{x}$ as oriented in the positive $x$-axis direction. Our findings revealed a consistent trend: both force [Fig. 3(a)] and torque [Fig. 3(b)] initially increase with an increment in drift velocity ($v_{d}$), irrespective of the particle’s material. However, at higher $v_{d}$ values, these quantities tend to decrease. This trend is consistent with the energy transfer.
@@
Figure 4: (a) Net power and (b) force for the NaCl particle as a function of $v_{d}$ at various graphene temperatures $T_{e}$, with $T_{p}=300\,$K and $d=0.2\,\mu$m. The power transfer curve at $T_{e}=300\,$K is multiplied by $10^{3}$ for better visibility. (c) Net power and (d) force versus $v_{d}$ with $T_{p}=300\,$K and $T_{e}=320\,$K. Its characteristic shape follows from the collective contributions of forward ($q_{x}{>}0$) and backward ($q_{x}{<}0$) SPPs, which are plotted separately as dashed lines.
@@
To gain a deeper understanding of the force and torque dynamics acting on the particle, we analyze the exchange of linear momentum and spin angular momentum between the forward ($q_{x}{>}0$) and backward ($q_{x}{<}0$) SPPs and the particle. When photons transfer from the graphene to the particle ($q_{x}{>}0$), the particle gains momentum, contributing to the lateral force along the positive $x$ direction. Conversely, when $q_{x}{<}0$, the particle loses momentum due to the reverse photon transfer, producing a lateral force in the same direction. Therefore, SPPs with both positive and negative $q_{x}$ generate a net lateral force directed along the positive $x$ direction. Due to the spin-momentum lock of SPPs Bliokh et al. (2015), the transverse spin of SPPs strongly couples with their propagation direction. As a result, when SPPs propagate in the positive (negative) $x$ direction, their spin aligns with the negative (positive) $y$ direction, respectively. This spin angular momentum transfer is facilitated by photon tunneling from the graphene to the particle in the case of $q_{x}{>}0$, generating a torque directed along the negative $y$ direction. Importantly, the torque stemming from $q_{x}{<}0$ contributions also points in the same negative $y$ direction. Consequently, polaritons in both positive and negative $q_{x}$ directions collectively generate a net lateral torque along the negative $y$ direction. The consistency of these findings is supported by Fig. 3(c), which depicts the torque spectral density $m_{y}$ as a function of $q_{x}$ and $q_{y}$ at $\hbar\omega=27.8\,$meV and $v_{d}=0.3v_{F}$. Notably, for both positive and negative $q_{x}$, the torque spectral density $m_{y}$ consistently maintains positive values, aligning closely with the dispersion curve of graphene plasmons. Remarkably, the force and torque induced by current-biased graphene are on par with those in thermally excited nonreciprocal surface electromagnetic waves Khandekar et al. (2021).
@@
Finally, we consider the scenario where the temperature of the particle differs from that of the graphene. Figure 4(a-b) shows the dependence of net energy transfer and lateral force on the NaCl particle with a radius of $R=40\,$nm at various temperatures of the graphene, while the temperature of the particle remains at 300300300\,300K. Interestingly, the energy transfer from the graphene to the NaCl particle increases by three orders of magnitude. To explain this, we calculate both the forward and backward SPPs that contribute to the energy and momentum transfer from the graphene to the particle in the case of $T_{e}{>}T_{p}$. The numerical results are presented in Fig. 4(c,d). From Fig. 4(c), we observe that the energy transfer for photons with opposite momentum has the same sign, which is completely different from the case when the temperatures are equal. Therefore, the energy transfer is greatly enhanced when a temperature difference is introduced. In contrast, we observe that the photons with $q_{x}{>}0$ and $q_{x}{<}0$ make positive and negative contributions to the force, respectively. By summing up the contributions from all photons, we find that the force experienced by the particle does not differ significantly from that in the equal temperature case. As a result, the temperature difference has minimal impact on the lateral force while significantly enhancing the net power transfer between the graphene and the particle.

### 16 Feb 2024 v3

@@ Abstract
In this paper, we study the near-field radiative energy, linear-momentum, and angular-momentum transfer from a current-biased graphene to nanoparticles. The electric current through the graphene sheet induces nonequilibrium fluctuations, causing energy and momentum transfer even in the absence of a temperature difference. The inherent spin-momentum locking of graphene surface plasmons leads to an in-plane torque perpendicular to the direction of the electric current. In the presence of a temperature difference, the energy transfer is greatly enhanced while the lateral force and torque remain within the same order. Our work explores the potential of utilizing current-biased graphene to manipulate nanoparticles.
@@ I Introduction
Nonreciprocity-induced lateral force has been studied recently Shapiro (2017); Gelbwaser-Klimovsky et al. (2021); Khandekar et al. (2021); Tsurimaki et al. (2023). An atom or nanoparticle placed above a sample with mobile carrier drift experiences a fluctuation force Shapiro (2017). In a system where vacuum-separated plates are at different temperatures, if at least one plate is nonreciprocal material, the setup can function as a heat engine where thermal energy is converted into mechanical work Gelbwaser-Klimovsky et al. (2021). When a dipolar particle is placed in the vicinity of nonreciprocal plasmonic slab, it experiences a lateral force and torque in the presence of a temperature difference between the particle and the slab Khandekar et al. (2021). The force is linked to the nonreciprocity of the surface polaritons and the torque to the spin-momentum locking of the surface polaritons.
@@
Current-biased graphene is also known for its nonreciprocity Borgnia et al. (2015); Duppen et al. (2016); Morgado and Silveirinha (2018, 2020). The exceptional electron mobility of graphene allows it to sustain nonreciprocal surface plasmon polaritons (SPPs) when driven by an electric current, which gives rise to interesting properties such as Fizeau drag Dong et al. (2021); Zhao et al. (2021) and negative Landau damping Morgado and Silveirinha (2017). Current-biased graphene is in a nonequilibrium state and this leads to a finite photonic chemical potential Henry and Kazarinov (1996) that depends on the in-plane wave vector for the thermal electromagnetic radiation. The occupation number of radiative photons is nonreciprocal, that is asymmetric with respect to the positive and negative in-plane wave vectors Volokitin and Persson (2008, 2011a, 2011b); Peng and Wang . The nonreciprocal properties of current-biased graphene can be adjusted through chemical doping or gate voltage, making it suitable for a wide range of applications in controlling thermal radiation. For example, near-field radiative heat transfer has been studied between two suspended graphene sheets Zhang et al. (2020a), graphene covered substrates Shi et al. (2022), graphene-nanoparticles Zhang et al. (2020b), and graphene-based multilayer systems Zhou et al. (2020). However, the physical aspects of the force and torque experienced by the nanoparticle near the current-biased graphene remain unexplored to date.
@@
In this paper, we investigate the near-field transfer of energy, linear momentum, and angular momentum between an isotropic dipolar nanoparticle and the current-biased graphene. The nonequilibrium state in the current-biased graphene induces energy transfer between the nanoparticle and the graphene even in the absence of a temperature difference. Importantly, these SPPs carry both linear and spin-locked angular momentum. As a result, there is an exchange of net linear and spin angular momentum between the particle and the SPPs, leading to both lateral force and torque experienced by the nanoparticle.
@@
Figure 1: (a) Schematic plot of the lateral force and torque exerted on an isotropic dipolar particle above a current-biased graphene in the near-field regime. The drift velocity $v_{d}$ is along the $x$ direction. (b) Dispersion of the SPPs of the current-biased graphene. The electric current effectively drags the SPPs which leads to the property of nonreciprocal propagation. (c) The photon occupation number difference $\delta n(\omega,q_{x})$ vs $\omega$ with $v_{d}=0.7v_{F}$.
@@ II System and formalism
The system is illustrated schematically in Fig. 1(a), where the spherical nanoparticle with radius $R$ is placed near the surface of current-biased monolayer graphene. Without losing generality, we assume that the drift velocity $v_{d}$ in graphene is along the positive direction of the $x$ axis. Within the framework of fluctuational electrodynamics, the total power $H$, lateral force along the $x$ direction $F_{x}$, and torque along the $y$ direction $M_{y}$ can be expressed as
$P=\int_{0}^{\infty}\frac{d\omega}{2\pi}p(\omega),$ (1)
with $P\in\{H,F_{x},M_{y}\}$ and the corresponding spectral densities $p\in\{h,f_{x},m_{y}\}$. The spectral densities are given by Khandekar et al. (2021)
$h(\omega)$ $=-\hbar\omega k_{0}^{2}{\rm Im}[\alpha(\omega)]\int\frac{d^{2}{\bm{q}}}{(2\pi)^{2}}\Theta(\omega,{\bm{q}})\delta n(\omega,q_{x}),$ (2)
$f_{x}(\omega)$ $=\hbar k_{0}^{2}{\rm Im}[\alpha(\omega)]\int\frac{d^{2}{\bm{q}}}{(2\pi)^{2}}q_{x}\Theta(\omega,{\bm{q}})\delta n(\omega,q_{x}),$ (3)
$m_{y}(\omega)$ $=-\hbar{\rm Im}[\alpha(\omega)]\int\frac{d^{2}{\bm{q}}}{(2\pi)^{2}}q_{x}{\rm Im}\big{(}r_{p}e^{2i\beta_{0}d}\big{)}\delta n(\omega,q_{x}),$ (4)
@@
$\Theta(\omega,{\bm{q}})={\rm Re}\bigg{[}\frac{r_{p}e^{2i\beta_{0}d}}{2\beta_{0}}\left(\frac{2q^{2}}{k_{0}^{2}}-1\right)\bigg{]},$ (5)
where the Fresnel reflection coefficient for $p$ polarization is given by
$r_{p}=\frac{\beta_{0}\sigma_{g}}{2\epsilon_{0}\omega+\beta_{0}\sigma_{g}}.$ (6)
Here, $\sigma_{g}$ is the surface conductivity of graphene, and $\epsilon_{0}$ is the dielectric constant in vacuum. The in-plane wave vector and the angular frequency are denoted by ${\bm{q}}=(q_{x},q_{y})$ and $\omega$. The magnitude of the out-of-plane wave vector in air is $\beta_{0}=\sqrt{k_{0}^{2}-q^{2}}$ with $k_{0}=\omega/c$ and $q=|{\bm{q}}|$. When an electric current is applied to the graphene so that the electrons near the Fermi surface experience a drift velocity, the system is not in local thermal equilibrium and the occupation number of the photons radiated from graphene is Doppler shifted. By assuming a drift velocity $v_{d}$ in the $x$ direction, the photon-occupation difference between the graphene and the nanoparticle is Volokitin and Persson (2008, 2011a, 2011b); Peng and Wang ; Tang et al. (2021b)
@@
where $T_{e}$ and $T_{p}$ are the temperatures of graphene and the nanoparticle, respectively. In getting the occupation number of the photons radiated from the current-biased graphene, one needs to apply the condition that $\hbar q_{x}v_{d}$ is much smaller than the chemical potential Peng and Wang . Equation (7) indicates that the electric current-induced nonequilibrium fluctuations can produce a finite energy flux even in the absence of a temperature difference. In the following, we assume that the temperature of the particle and the graphene is the same with $T_{p}=T_{e}=T=300\,$K, unless there is a special account for the temperatures. The energy transmission function of the energy flux in Eq. (2) is defined through $h(\omega)=\int d^{2}{\bm{q}}{\cal Z}_{h}(\omega,{\bm{q}})/4\pi^{2}$ with
${\cal Z}_{h}(\omega,{\bm{q}})=-\hbar\omega k_{0}^{2}{\rm Im}[\alpha(\omega)]\Theta(\omega,{\bm{q}})\delta n(\omega,q_{x}),$ (8)
which provides the information of the energy transfer at given $\omega$ and $\bm{q}$. Similar definitions are, respectively, applied to the force and torque.
@@
The dispersion relation of SPPs in a monolayer graphene can be obtained through Gonçalves and Peres (2016)
$2\varepsilon_{0}\omega+\beta_{0}\sigma_{g}=0,$ (9)
where the nonlocal surface conductivity $\sigma_{g}$ is given by Svintsov and Ryzhii (2019)
@@
with $\mu(T)=2k_{B}T\ln[2\cosh(\mu_{g}/2k_{B}T)]$. Here, $T$ is the temperature, $\mu_{g}$ is the chemical potential, $\gamma$ the damping parameter, and $e_{0}$ the electron charge. The nonlocal surface conductivity is obtained under the long-wavelength limit and by neglecting the interband transition which is justifiable for $2\mu_{g}>\hbar\omega$. In the numerical calculation, we set the Fermi velocity $v_{F}=10^{6}\,$m/s, the chemical potential $\mu_{g}=0.1\,$eV and the damping parameter $\gamma=3.7\,$meV. Figure 1(b) shows the nonreciprocal dispersions of SPPs in the presence of an electric current.
@@ III Numerical Results
Before considering the nonequilibrium energy and momentum transfer, we discuss the nonreciprocal properties of $\delta n(\omega,q_{x})$ shown in Fig. 1(c). We can see that the photon occupation number difference $\delta n(\omega,q_{x})$ is positive for positive $q_{x}$ under $\omega{>}q_{x}v_{d}$. This indicates that the effective photonic temperature of the graphene is higher than the particle temperature $T_{p}$. Furthermore, it should be noted that contributions from $\omega{<}q_{x}v_{d}$ can be disregarded. This is because these contributions become significant only when the separation between the particle and the graphene reaches subnanometer scales Volokitin and Persson (2011b), whereas in our case, the distance is on the order of micrometers. Consequently, photons, especially those with positive $q_{x}$, flow from the graphene to the particle. In contrast, for negative $q_{x}$, $\delta n(\omega,q_{x})$ is always negative. Therefore, the effective photonic temperature of the graphene is lower than $T_{p}$. This sign reversal indicates a reversal in the direction of photon transfer compared to the behavior exhibited by positive $q_{x}$.
We consider three kinds of nanoparticle: NaCl, undoped GaAs, and InSb. The intricate dielectric characteristics are given by the Lorentz-Drude model with
@@
where $\varepsilon_{\infty}$ is the high-frequency dielectric constant, and $\omega_{L}$ is the longitudinal optic-phonon frequency, $\omega_{T}$ is the transverse optic-phonon frequency, $\Gamma$ is the optic-phonon damping constant. Within the electrostatic limit, we can reasonably neglect multiple scattering interactions between the particle and the graphene, and describe the particle effectively using its electrical polarizability which is expressed as $\alpha(\omega)=4\pi R^{3}[\varepsilon(\omega)-1)]/[\varepsilon(\omega)+2]$ Zhao et al. (2012). It is important to note that these assumptions are valid only under the condition of $R\ll d\ll\hbar c/k_{B}T$. In Fig. 2(a), we present the imaginary part of the polarizability for these nanoparticles, revealing distinct dipolar resonances.
Now we study the effect of drift velocity on the net power transfer experienced by different nanoparticles, as shown in Fig. 2(b). We differentiate between positive and negative energy fluxes, which represent the flow of energy from the graphene to the particle and from the particle to the graphene, respectively. Our analysis reveals intriguing behaviors. For NaCl and InSb particles, the power transfer initially increases as $v_{d}$ increases. However, with increasing $v_{d}$ further, the power transfer decreases. In contrast, the direction of net power transfer undergoes a change for GaAs particles as the drift velocity increases. These are due to the interplay between the nonreciprocity of the SPPs and photon occupation number of graphene.
@@
To analyze the contributions of forward ($q_{x}{>}0$) and backward ($q_{x}{<}0$) SPPs to net power transfer, we use the NaCl particle as an example. In Fig. 2(c), we show the energy transmission function and the dispersion of graphene plasmons (shown by the cyan dashed line). This visual representation aligns the bright branches of the energy transmission function with the cyan dashed line, confirming the dominant role played by SPPs in governing power transfer. For $q_{x}{>}0$, the energy transfer coefficients are positive, indicating energy transfer from graphene to the particle. For $q_{x}{<}0$, the energy transfer coefficients are negative, signifying energy transfer from the particle back to the graphene. The net energy transfer is the sum of contributions from $q_{x}{>}0$ and $q_{x}{<}0$. At lower drift velocities, the nonreciprocal characteristics of SPPs are not pronounced, and energy transfer is mainly governed by the nonreciprocal photon occupation number difference $\delta n(\omega,q_{x})$, as shown in Fig. 1(c). However, as the drift velocity increases, nonreciprocal SPPs interplay with $\delta n(\omega,q_{x})$. The high drift velocities suppress the net power transfer due to the nonreciprocal photon occupation number even change the energy transfer direction between GaAs and graphene.
@@
In the following, we investigate the transfer of linear and angular momenta between the nanoparticles and the current-biased graphene. We define positive $M_{y}$ along the negative $y$ direction and positive $F_{x}$ as oriented in the positive $x$ direction. Our findings revealed a nonmonotonic behavior: Both force [Fig. 3(a)] and torque [Fig. 3(b)] increase with increasing the drift velocity up to a certain value, then start to decrease. This behavior is consistent with the energy transfer.
@@
Figure 4: (a) Net power and (b) net force for the NaCl particle as a function of $v_{d}$ at various temperatures $T_{e}$ in graphene with $T_{p}=300\,$K and $d=0.2\,\mu$m. The curve of power transfer at $T_{e}=300\,$K is multiplied by $10^{3}$ for better visibility. (c) Net power and (d) net force vs $v_{d}$ with $T_{e}=320\,$K and $T_{p}=300\,$K. The contributions from the forward ($q_{x}{>}0$) and backward ($q_{x}{<}0$) SPPs are separately shown.
@@
To have a deeper understanding of the force and torque acting on the particle, we analyze the exchange of linear momentum and spin angular momentum between the forward ($q_{x}{>}0$) and backward ($q_{x}{<}0$) SPPs and the nanoparticle. When photons transfer from the graphene to the particle ($q_{x}{>}0$), the particle gains momentum, contributing to the lateral force along the positive $x$ direction. Conversely, when $q_{x}{<}0$, the particle loses momentum due to the reverse photon transfer, producing a lateral force in the same direction. Therefore, SPPs with both positive and negative $q_{x}$ generate a net lateral force directed along the positive $x$ direction. Due to the spin-momentum locking of SPPs Bliokh et al. (2015), the transverse spin of SPPs strongly couples with their propagation direction. As a result, when SPPs propagate in the positive (negative) $x$ direction, their spin aligns with the negative (positive) $y$ direction, respectively. This spin angular momentum transfer is facilitated by photon transfer from the graphene to the particle in the case of $q_{x}{>}0$, generating a torque directed along the negative $y$ direction. Importantly, the torque contributed from $q_{x}{<}0$ is also along the negative $y$ direction. Consequently, the net lateral torque is along the negative $y$ direction. These are supported by Fig. 3(c) which shows the torque spectral density $m_{y}$ as a function of $q_{x}$ and $q_{y}$ at $\hbar\omega=27.8\,$meV and $v_{d}=0.3v_{F}$. Notably, for both positive and negative $q_{x}$, the torque spectral densities $m_{y}$ are positive. Remarkably, the force and torque induced by current-biased graphene are of the same order of magnitude as those generated by thermally excited nonreciprocal surface electromagnetic waves Khandekar et al. (2021).
@@
Finally, we consider the scenario where the temperature of the particle differs from that of the graphene. Figures 4(a) and  4(b) show the dependence of net energy transfer and lateral force on the NaCl particle at various temperatures of the graphene. The nanoparticle temperature is at 300300300\,300K. Notably, the energy transfer from the graphene to the nanoparticle increases by three orders of magnitude. To explain this, we calculate the contributions from both the forward and backward SPPs at $T_{e}{>}T_{p}$ [see Figs. 4(c) and (d)]. It can be seen from Fig. 4(c) that the energy transfer for photons with opposite momentum has the same sign, which is completely different from the case in the absence of a temperature difference. This means that the contribution from the temperature difference dominates over that from the electric current. Therefore, the energy transfer is greatly enhanced in the presence of a temperature difference. In contrast, the photons with $q_{x}{>}0$ and $q_{x}{<}0$ make positive and negative contributions to the force, respectively. We find that the force experienced by the particle does not differ significantly from that without a temperature difference. The torque has similar behavior as the force. Thus, the temperature difference has a marginal impact on the lateral force and torque while significantly enhancing the net power transfer with $T_{e}{>}T_{p}$.